OpenStudy (anonymous):

Solve x2 + 10x = -21. x = 7 and x = 3 x = -7 and x = -3 x = -7 and x = 3 x = 7 and x = -3

OpenStudy (anonymous):

First we need to put the -21 on the other side of the equal sign. So the equation reads\[x^2+10x+21\]

OpenStudy (anonymous):

Do you want to do the quadratic formula or factoring

OpenStudy (anonymous):

I'm not sure. Whatever one will work.

OpenStudy (anonymous):

Both will work, I'll do it both ways i guess. \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] a=1 b=10 c=21 \[\frac{ -10 \pm \sqrt{10^2-4(1)(21)} }{ 2(1) }\] \[\frac{ -10\pm \sqrt{100-84} }{ 2 }\] \[\frac{ -10 \pm \sqrt{16}}{ 2 }\] This will break into 2 equations \[\frac{ -10 + 4 }{ 2 }\] and \[\frac{ -10-4 }{ 2 }\]

OpenStudy (anonymous):

Can you solve the last 2 for me plz?

OpenStudy (anonymous):

-3 and -3 ?

OpenStudy (anonymous):

you got the first one right not the second one tho (a negative minus a number will decrease the number) \[\frac{ -14 }{ 2 }\]

OpenStudy (anonymous):

-7? @KlOwNlOvE

OpenStudy (anonymous):

sorry was afk, yes the solution set would be -7 -3